OpenStudy (anonymous):

The smallest angle in a traingle is half the largest angle. The middle angle measures 30* less than the largest angle . Find the measure of each angle.im

OpenStudy (anonymous):

let x,y.z be the angles

OpenStudy (anonymous):

x is smallest , y is middle , z is largest

OpenStudy (anonymous):

x = z/2

OpenStudy (anonymous):

y = z - 30

OpenStudy (anonymous):

why is the z/2?

OpenStudy (anonymous):

becuse it is given that smallest angle is half the largest angle

OpenStudy (anonymous):

x is smallest and z is largest

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now as x+y+z = 180 z/2 + (z-30) + z = 180

OpenStudy (anonymous):

z/2 +2z = 210

OpenStudy (anonymous):

5z/2 = 210 z = 84

OpenStudy (anonymous):

x = z/2 84/2 42 y = z-30 = 84-30 = 54

OpenStudy (anonymous):

so configuring al this what is the measure of each?

OpenStudy (bahrom7893):

84, 42 and 54 apparently

OpenStudy (anonymous):

oh theres 3 of them?

OpenStudy (bahrom7893):

It's a triangle of course there are 3 angles..